OpenStudy (anonymous):

Determine the value of x. DO,k(x, 9) → (–10, –6) A. 5 B. 7 C. 12 D. 15

OpenStudy (anonymous):

translation rule is x-15 so if x-15=-10 =>x=-10+15 =>x=5 optioon A...

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

k (x, 9) = (-10, -6) that means (k*x, k*9) =(-10,-6) Which gives you k*x = -10 k*9 =-6 From them, find k from the second one, plug that k into the first one to get x

OpenStudy (anonymous):

Show your attempt, or just the answer, PLEASE, don't guess.

OpenStudy (anonymous):

thank you